OpenStudy (anonymous):

In October, the average price of aluminum was 87 2/3 cents per pound. during that time Severo recieved 526 cents for aluminum cans that he sold for recyclingat a scrap metal center. Assuming that he recieved the average price, determine how many pounds of aluminum cans severo recycled. Type a whole number, proper fraction,or mixed number.

OpenStudy (shane_b):

All of the problems you're posting are essentially the same. This one would be:\[\frac{526}{87\frac{2}{3}}=?\]

OpenStudy (anonymous):

how do I solve this problem.

OpenStudy (anonymous):

do i write that in a improper fraction

OpenStudy (anonymous):

88

OpenStudy (anonymous):

526 divided by 88

OpenStudy (shane_b):

Yes!

OpenStudy (anonymous):

i got 5.97 what is the fraction of that.. can you please explain because this where i get confuse at

OpenStudy (shane_b):

Ok...let's do the whole thing. If you do it without converting to a decimal you won't be left with a fraction.

OpenStudy (anonymous):

would it be 5/9

OpenStudy (shane_b):

\[\frac{526}{87\frac{2}{3}}\] First convert 87 2/3 into an improper fraction:\[\frac{87*3+2}{3}=\frac{263}{3}\]Now do the division:\[\frac{526}{\frac{263}{3}}\]\[526*\frac{3}{263}=\frac{1578}{263}=6\]

OpenStudy (shane_b):

So he recycled 6 pounds of aluminum.

OpenStudy (anonymous):

ok. I am just so confuse with alegbra. its complicatedd for me. I sure do appreciate how you explain everything too me. thanks you so much.

OpenStudy (anonymous):

can you help me with 2 more problems

OpenStudy (shane_b):

I have to go but I'm sure someone else can help you